I am using Cr (Chromium) metal, which forms Cr3+ ions when dissolved. If 0.134 g of the metal dissolve, how moles of electrons passed through the circuit?
1
Expert's answer
2012-09-06T11:53:03-0400
Cr 0 -----> Cr3+ + 3e

n = 0.134/51.99=0.002577
moles

0.002577*3=0.007732 moles